Test the consistency of the following system x+6y=9, 3x+18y=6.


1
Expert's answer
2021-03-31T15:25:30-0400

The equations are


"x+6y=9\\
3x+18y=6"


Multiplying the first equation by 3 we get 3x+18y=27


From second equation we get 3x+18y=6. Therefore 6=27 from last paragraph, which is not true.

Hence the system of equation is inconsistent.